OpenStudy (anonymous):

Write the equation of a line in slope-intercept form that has a slope of -2 and passes through the point (2, -8).

OpenStudy (anonymous):

Use the point slope equation. y= mx + b -8 = -2(2) + b -8 = -4 + b -4 = b So: y=-2x -4